OpenStudy (anonymous):

Please help: Solve 1/3x = -4/5

OpenStudy (anonymous):

Cross multiply

OpenStudy (anonymous):

So multiply 3x times -4 and 1 times 5 That would leave you with -12x = 5 Can you solve from there?

OpenStudy (anonymous):

is the x above or below the fraction eg, 1x/3, (1/3)x

OpenStudy (anonymous):

\[\frac{ 1 }{ 3x } = \frac{ -4 }{ 5 }\]

OpenStudy (superdavesuper):

I could be wrong but I thought the Q is \[\frac{ 1 }{ 3 }x=-\frac{ 4 }{ 5 }\]

OpenStudy (anonymous):

thats exactly what i thought, that is why i was asking for clarification

OpenStudy (anonymous):

sorry i am back. the x variable is being multiplied

OpenStudy (anonymous):

could be 4/15, -4/15, -2 2/5, or 2 2/5

OpenStudy (anonymous):

Dave's example is how it looks

OpenStudy (superdavesuper):

so x=-4/5*3

OpenStudy (anonymous):

i thought that was it, many thanks :)